CRY+VIS Interface Meeting
Date: 2021/12/17 13:00-
- IX: Finalization. Closing of the chamber on 12/14. Adjustment of the BFV. Glich in the F3 signals.
- EX: Finalization. Closing of the chamber on 12/15. Pump down on 12/16.
- EY: There is a problem (BF touching).
- Replacement of flip cables:
- TM height setting: ITMY.
- Oplev: Setup of the TM Oplev in IY after the TM alignment.
- ACC: Assembly for EY is under way.
- Commissioning in air: Return of GRY.
- Commissioning in vacuum: Pump down of EXC/EXV on 12/15.
- Overall commissioning:
- Acceptance check: Review on 12/10.
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers: TMP to IXV and EXC-BRT on 12/13.
- Heater control:
- Power source for Oplev:
Date: 2021/12/17 14:00-
- Adjusted the height of the BF damper in IX.
- Investigating the glitches in the F3 signals in IX.
- Monitored the displacement of the Type-A tower during the pump down in EX.
- Concerning the BS F1 LVDT calibration, we tested the hardware to hold the displacement sensor above the BF. It seems to work well, although it was difficult to install it. We were not expected it to be easy, so it's fine. In principle, we aim to do the calibration next week, although that would depend on other tasks we have to do.
- The temperature became closer to O3GK temperature around the BS.
- We removed 5.7 kg of ballast mass from the BS IP:
- The resonant frequencies changed
L: 44 ==> 78 mHz.
T: 79 ==> 97 mHz.
Before removing mass, the resonant frequency in L seemed to depend on the IP position. We didn't check after removing mass, but we should.
- The resonant frequencies changed
- Concerning the wandering peak noise in BS F1 LVDT, we prepared to measure the voltage of the BS chamber with respect to the electrical ground. We were about to connect the cables and we were distracted by the PR2 issue. We aim to do this next week.
- Terrence modified SR2 Guardian on Friday last week, but this hadn't been reported in this meeting.
- While adjusting the yaw of the suspension with the traverser, the suspension suddently moved a lot in yaw and in pitch also, and the traverser jammed.
- Mechanical problems detected by visual inspection:
- At least two traverser motors reached their mechanical limits.
- Two magnets from the mirror fell off out of the coil body cavity, which was surprising.
- A security structure stud fell off from the ring below the BF onto the suspended breadboard. Clearly, the stud had not been fastened correctly.
- Judging from the position of at least one of the BF damper primary coils, the BF had a lot of tilt.
- On top of the BF there was a cable that was hooking to one of the picomotor assemblies. We don't remember seeing this before, so it migth have been caused by the suspension moving too much in yaw.
- Apparently, a human error caused the problem.
- According to Ikeda-san, the suitable way for us to operate the motors is to override the position of the motor that is stored in the driver.
- He might have forgotten to do this, with the consequence of the driver using an arbitrary value for the current position, thus moving the motor too much.
Another concern would be the velocity, but I didn't find information about this in Ikeda-san's klog. I'll ask him.
- The traverser has some switches, but they are not limit switches, they are for defining the origin of the coordinate system used to measure the position of the stepper motors.
- As described in the klog:
- We locked the suspension.
- Unhook the cable.
- Moved the stepper motors by hand (surprisingly, this was easy to do) until they were centered.
Released the suspension. (We forgot to check the BF tilt by visual inspection after releasing it.)
- Checked that the green laser reached the X End. The alignment of the green laser depends on the alignment of the PR2 mirror to a certain extent due to its small radius of curvature.
- We closed the chamber.
- The recovery work has begun.
- Sato-san and Hirata-san began removing the adhesive from the magnet bases.
Fortunately, it seems that all the adhesive remained on the magnet bases and not on the mirror surface.
- Aso-san will buy fresh EP-30. It has 6 months of shelf life.
- Takahashi-san searched for the jigs to glue the magnets back onto the mirror.
- The RM is not designed to be a jig for the mirror, so careful alignment of the mirror with respect top the RM is required.
- This alignment is done by visual inspection in pitch, yaw and transverse (in the vertical direction also, but this direction is constrained by the wire length).
- It seems impossible to do this type of work with the mid-size baffle in the way.
- Type-A team: BF in EY
- Type-B team: Recovery of PR2
- The helmet dropped from the second floor in the PR2 booth.
- Part of a clamp of a floor panel fell from the second floor next to the BS frame. It seems there are a few loose clamps.
(Date in Kamioka)
The week from 12/20
(F2F meeting in 12/20-21)
- Takahashi: 12/20-24
- Sato: 12/20-24
- Hirata: 12/20-24
The week from 12/27
(Holidays in 12/29-1/3)